# Model for Peanuts: Hijacking ML models without Training Access is Possible

### v1

@@ 7.1. Context
This section illustrates our attack with a use case in the healthcare domain. Let the original task $\mathcal{T}$ be Pulmonary Disease Diagnosis (PDD) using chest X-ray images. The original task is a binary classification of whether the patient has ’Pneumonia’. Such a model can be deployed within MLaaS API such as Google Healthcare API (Google, [n. d.]). Patients or practitioners can query the model with X-ray images and receive feedback on potential pneumonia diagnoses. Our attack aims to investigate the possibility of inferring even more information on the type of disease using the same model. Specifically, the attacker tries to infer a hijacking task on samples identified with pneumonia, i.e., whether the individual’s infection is from a viral or bacterial origin. Furthermore, we assume that the adversary has access to a database of X-ray images labeled with the type of infection. These images, which may be publicly available, are not necessarily associated with the individuals querying the PDD model.
7.2. Setup
@@
Dataset: Chest X-Ray Images (Pneumonia) This dataset (Kermany et al., 2018) consists of $5,863$ X-ray images categorized into ’Pneumonia’ and ’Normal’ classes. Two expert physicians graded the diagnoses, and a third expert validated the evaluation set to ensure diagnostic accuracy. The ’Pneumonia’ category is further labeled into ’Viral’ and ’Bacterial’ subcategories.
We consider the same backbone architectures as in Section 5.2 with slight modifications to enable the (original) binary classification.
7.3. PDD Models recognize more than Pneumonia
The adversary aims to hijack a PDD model to extract more information on the victim’s health. Specifically, from unknown query X-ray images only using the PDD model’s output, the attacker aims to infer the pulmonary infection type (i.e., viral or bacterial). Such information can be highly sensitive and might be further used for malicious intent. In the rest of this section, we consider the classification of pulmonary infection to be our hijacking task.
Following our threat model, we conduct a hijacking attack using BEK from the hijacking dataset to recognize the infection type of unknown query X-ray images. Similar to our evaluation of SnatchML  effectiveness in Section 5.3, we use the X-ray images from the test dataset as targets for the hijacking attack.
Figure 8. PDD Attack: Accuracy of the original task and hijacking attacks. Random guess probability is at 0.5.
Figure 9. Results of SnatchML  on MobileNet pretrained for image classification on ImageNet and Cifar10. The models are hijacked to infer other unrelated tasks defined by their datasets on the x-axis.
@@ 7.3. PDD Models recognize more than Pneumonia
Figure 8 reports the accuracy results of our hijacking attacks under different settings and models. We observe that the PDD model achieves an accuracy up to $\sim 85\%$ on the original task. Interestingly, the accuracy of the hijacking task (i.e., type of pneumonia infection) under white-box settings (using FV) is comparable to the accuracy of the model in the original task. Among all architectures, SnatchMLwas particularly successful on ResNet-9 under both white and black-box settings, reaching an attack accuracy of $76\%$ and $78\%$, respectively. Despite its inherited problem of learning complex tasks on small datasets, the Transformer achieves a hijacking accuracy of $62.5\%$. The disparity in our attack efficiency as a function of the model architecture for the same settings provides an interesting insight into the importance of the model hyperparameters in the underlying phenomena leading to the hijacking success. It is also worth noticing that under the white-box setting and for all models, there’s a strong correlation between the accuracy of the original and hijacking tasks. This further emphasizes the overlap in the learning dynamics for the original and hijacking tasks. Such an overlap can be of a high risk of privacy breaches, especially for sensitive applications like healthcare systems. SnatchMLachieves results that are close to the hijacking UB.
Figure 10. Correlation between the size of benign models trained on original task (ER on CK+ dataset) and their effectiveness in learning hijacking tasks: The results depict the variations in the accuracy of the hijacking tasks for user re-identification (First row) and user identification (Second row) when scaling the ML model size via model’s width expansion. The hijacked models are of different architectures (per column): 2D-CNN, ResNet-9, MobileNet, and Transformer.
8. What if the hijacking task is unrelated to the original task?
@@
Previous experiments demonstrate that a model trained for an original task can be effectively hijacked to perform a different task. This finding has so far involved scenarios where the hijacking tasks share semantic relatedness or overlap in data distribution with the original task. While this represents a significant security threat, the requirement for task-relatedness can be viewed as a limitation from an attack perspective. State-of-the-art model hijacking attacks are more general and are not particularly constrained by the relatedness to the original task. In this section, we extend our investigation to examine the generalizability of SnatchML  to hijacking tasks unrelated to the original task.
@@
Experimental setup. We evaluate several image classification benchmarks as original tasks. We use a MobileNet pretrained on ImageNet and CIFAR-10, sourced from the Pytorch model Zoo (Pytorch, [n. d.]). These pretrained models are the targets of SnatchML, which aims to hijack them for unrelated tasks involving different data distributions. Specifically, we hijack the models for MNIST, CK+ for ER, CK+ for biometric identification, and Olivetti for face recognition. We run SnatchML  on the test distribution to evaluate the hijacking task performance. For each hijacking attack, Figure 9 presents the performance under both black-box and white-box settings and the probability of a random guess.
@@
Observations. The results in Figure 9 illustrate the capacity of overparametrized models to extract useful knowledge from a data distribution unrelated to their original training dataset. In fact, the SnatchML  hijacking task accuracy is comparable to the state-of-the-art accuracy of tasks like MNIST when the victim model is trained on ImageNet. In the following section, we attempt to provide possible explanations for these results.
9. Why do ML models learn more than they should?
@@
In the previous sections, we have shown that SnatchML can hijack ML models to perform tasks that can be related or entirely unrelated to the original task. While the first case is intuitive, as illustrated in Section 2, the second case is more surprising. In the next two subsections, we explore the reasons behind these observations by examining two hypotheses: (i) Overparametrization of the ML model and (ii) Random Projection.
9.1. Training ML models do not respect the ”Least Privilege Principle”
@@
ML models may learn features that are not particularly related to the original task. For example, in the context of ER, these models might capture facial characteristics that could inadvertently be used for biometric identification. From a conceptual perspective, such a phenomenon might lead to critical security breaches due to not respecting the fundamental principle of ”least privilege” in cybersecurity. Practically, the amount of parameters in an ML model reflects the upper bound of its learning capacity. The overparametrization implies that such models contain more parameters than are necessary to accomplish the original task. In both software (Wang et al., 2013) and hardware (Fern et al., 2015) systems, undefined behavior and don’t-care states have been shown to be potential sources of vulnerabilities. In the same vein, we believe that this extra capacity represents undefined behavior and don’t-care states with respect to the model that can be exploited by attackers. We posit that the success of our hijacking attack is related to the overparametrization of victim models. We suggest that the model’s capacity to unintentionally learn unnecessary features positively correlates with the accuracy of the hijacking task. This is particularly supported by the fact that we’re exploiting the same extracted feature from the original model to perform hijacking tasks. We assume that within those features, some strongly correlate with the hijacking task (e.g., ER vs. biometric identification).
@@
To verify this hypothesis, we study the correlation between the ML model size and accuracy of the hijacking tasks for user re-identification and biometric identification. The results of our study are shown in Figure 10. We study the case of four ML models, notably 2D-CNN, ResNet-9, MobileNet, and Transformer. We vary the width expansion ratio (i.e., number of channels) from 0.25×0.25$\times$0.25 × to 2.5×2.5$\times$2.5 × to emulate different amounts of model parameters while maintaining the same baseline model architecture. From the reported results in Figure 10, we draw the following insights:
@@
Third, the observed trends between the model size and the learning capacity for the hijacking task differ w.r.t the task complexity. For instance, ResNet-9 shows slight improvement of $2\%$ in the hijacking attack for user re-identification under overparameterized regime. For the identification task in Olivetti, ResNet-9 depicts higher improvement of $9\%$ in the hijacking attack ($2^{nd}$ row, $1^{st}$ and $2^{nd}$ columns in Figure 10). The re-identification task (i.e., the ER training dataset) includes $89$ user identities, making the re-identification task much more challenging than the identification task, which involves around $40$ user identities.
Figure 11. Hijacking attack on a randomly initialized model to perform ER, Identification, as well as MNIST and CIFAR100. Rnd refers to random guess for each benchmark.
9.2. General case: Universality of Random DNNs
@@
The general case corresponds to a model hijacked for a task totally unrelated to the original task. The rationale by which features learned by the model on the original task overlap with the hijacking task does not hold under these assumptions. As far as the hijacking task is concerned, this case is similar to a randomly initialized model. We hypothesize that the universal approximation capabilities of random DNNs can explain the relatively high accuracy of the hijacking task. We recall the Johnson-Lindenstrauss Lemma (Johnson and Lindenstrauss, 1984), which states that a set of points in a high-dimensional space can be projected into a lower-dimensional space while approximately preserving the pairwise distances between the points. While this lemma might give an interesting insight, it cannot rigorously explain the behavior of random ML models because of their non-linearity. However, several works such as (Giryes et al., 2016; Basteri and Trevisan, 2023; Liao and Couillet, 2018; Rahimi and Recht, 2007) investigated the capabilities of random neural networks. Particularly, Giryes et al. (Giryes et al., 2016) prove that DNN preserves the metric structure of the data as it propagates along the layers. Interestingly, our work shows that networks tend to decrease the Euclidean distances between points with a small angle between them (“same class”) more than the distances between points with large angles between them (“different classes”). Further, the theoretical underpinnings provided by (Giryes et al., 2016) demonstrate that deep random networks can act as powerful feature extractors, where the depth and width of the network play crucial roles in enhancing the representational power of the features.
@@
While the properties of random neural networks are extensively studied in the ML community, our work shows the potential security threats that can emanate from these properties. To illustrate this perspective, we run our hijacking attack on totally random neural networks, i.e., the parameters are sampled from a random Gaussian distribution. The results shown in Figure 11 are coherent with our observations in Section 8. In the Appendix, we provide a visualization of this setting in the latent space in Figure 13.
10. Countermeasures
In this section, we investigate possible countermeasures to mitigate the threat of SnatchML; We propose the following two approaches:
10.1. Meta-unlearning: meta-Learning-based malicious task unlearning
In cases where the hijacking threat can be identified at training time with a precise task, e.g., biometric identification, we propose to train the model to learn the original task and simultaneously unlearn the potentially malicious one.

### v2

@@ VI Scenario 3: Pneumonia Diagnosis
This section illustrates our attack with a use-case in the healthcare domain, where the original task is Pulmonary Disease Diagnosis (PDD) using chest X-ray images, i.e., a binary classification of whether the patient has ’Pneumonia’. Such a model can be deployed within MLaaS API such as Google Healthcare API [19]. Patients or practitioners can query the model with X-ray images and receive feedback on potential pneumonia diagnoses. Our attack investigates the possibility of inferring more information on the type of disease using the same model. Specifically, the attacker tries to infer a hijacking task on samples identified with pneumonia to infer whether the individual’s infection is from a viral or bacterial origin. We assume that the adversary has access to a database of X-ray images labeled with the type of infection. These images, which may be publicly available, are not necessarily associated with the individuals querying the PDD model.
VI-A Setup
@@
Chest X-Ray Images (Pneumonia) Dataset. This dataset [25] consists of $5,863$ X-ray images categorized into ’Pneumonia’ and ’Normal’ classes. Two expert physicians graded the diagnoses, and a third expert validated the evaluation set to ensure diagnostic accuracy. The ’Pneumonia’ category is further labeled into ’Viral’ and ’Bacterial’ subcategories. We consider the same back-bone architectures as in Section IV-A with slight modifications to enable the (original) binary classification.
VI-B PDD Models recognize more than Pneumonia
@@
The adversary aims to hijack a PDD model to extract more information on the victim’s health record. Specifically, from unknown query X-ray samples, only using the PDD model’s output, the attacker aims to infer the pulmonary infection type (i.e., viral or bacterial). Figure 7 shows the accuracy results of SnatchML  under different settings. We observe that the PDD model achieves an accuracy up to $\sim 85\%$ on the original task. Interestingly, the accuracy of the hijacking task (i.e., type of pneumonia infection) is comparable to the accuracy of the model in the original task. Among all architectures, SnatchML was particularly successful on ResNet-9 under both white and black-box settings, reaching an attack accuracy of $76\%$ and $78\%$, respectively. Despite its inherited problem of learning complex tasks on small datasets, the Transformer achieves a hijacking accuracy of $62.5\%$. The disparity in attack efficiency as a function of the model architecture for the same settings provides an interesting insight on the importance of the model hyperparameters in the underlying phenomena leading to the hijacking success. Note that under white-box and for all models, there’s a strong correlation between the accuracy of the original and hijacking tasks. This further emphasizes the overlap in the learning dynamics for the original and hijacking tasks. SnatchML achieves results that are close to the hijacking UB despite being with no training access.
Figure 7: PDD Attack: Accuracy of the original task and hijacking attacks. Random guess probability is at 0.5.
VII Scenario 4: ECG Diagnosis
In this section, we evaluate our attack on a modality other than images. Specifically, we focus on ECG signals, which capture cardiac electrical activity over time and produce one-dimensional signal data. To assess the susceptibility of ECG signals to hijacking attacks at inference time, we consider the following settings:
Case (i): The original task $\mathcal{T}$ is ECG-based arrhythmia classification into $5$ categories: ’normal’ and $4$ anomalous categories. The adversary’s goal is to hijack a model pre-trained for arrhythmia detection to perform biometric identification, attempting to identify individuals from unknown ECG queries. These ECG signals are sourced from a dataset different from the one used for the original task.
Case (ii): The original task is a binary classification of normal versus arrhythmic signals. The adversary aims to repurpose the model to identify specific types of cardiac irregularities within the ECG data, categorizing them into $5$ predefined classes.
VII-A Setup
MIT-BIH Arrhythmia Dataset [32]: includes ECG recordings from $47$ participants. The dataset contains over $110,000$ samples, categorized into $5$ classes: N (Normal), S (Supraventricular ectopic beat), V (Ventricular ectopic beat), F (Fusion beat), and Q (Unknown beat).
ECG-ID [29]: comprises $310$ ECG recordings from $90$ individuals and labeled by identity. To process ECG 1D-inputs, we use $3$ models: 1D-CNN, 1D-ResNet-18, and 1D-Inception.
VII-B Hijacking arrhythmia classifiers for biometric identification
The results of the first scenario are depicted in the first row of Figure 8 across the different architectures (1D-CNN, 1D-Inception, and 1D-ResNet-18). For example, 1D-Inception ranges from $29.68\%$ (Top-1) to $54.19\%$ (Top-5) when the attack is black-box. These results illustrate the models’ effectiveness in biometric identification compared to a random accuracy of $0.11\%$. This underscores the models’ capability to generalize beyond their original task of ECG classification, posing notable implications for security and privacy.
Figure 8: SnatchML performance on hijacking ECG: 1st Row: biometric identification on ECG-ID dataset. 2nd Row: Binary classification of cardiac abnormalities on the MIT-BIH Arrhythmia dataset.
VII-C Binary Classification Models learn arrhythmia categories
Figure 8 demonstrates the vulnerability of the three models in a hijacking scenario with binary classification of ’normal’ versus ’anomalous’ ECG as original task. SnatchML  aims to hijack the models for ECG classification into five distinct categories of Arrhythmia (see the second row of Figure 8). The models achieve relatively high accuracy reaching $79\%$ 1D-ResNet-18, $59\%$ for 1D-Inception, and $60\%$ for 1D-CNN while the hijacking UB accuracy is up to $98\%$ This setting underscores the models’ ability to generalize across different types of cardiac irregularities, highlighting the hijacking risk in sensitive domains such as healthcare.
VIII What if the hijacking and original tasks are unrelated?
@@
Previous experiments demonstrate that a model trained for an original task can be effectively hijacked to perform a different task. This finding has so far involved scenarios where the hijacking tasks share semantic relatedness or an overlap in data distribution with the original task. While this represents a significant security threat, the requirement for task relatedness can be viewed as a limitation from an attack perspective. In fact, state-of-the-art model hijacking attacks are more general and are not particularly constrained by the relatedness to the original task. In this section, we extend our investigation to examine the generalizability of SnatchML to hijacking tasks that are unrelated to the original task.
Figure 9: Results of SnatchML on MobileNet pretrained on ImageNet and Cifar10. The models are hijacked to infer other unrelated tasks defined by their datasets on the x-axis.
VIII-A Unrelated Image Classification Tasks
@@
Setup. We evaluate several image classification benchmarks as original tasks. We use a MobileNet pretrained on ImageNet and on CIFAR-10, from the Pytorch model Zoo [36]. These pretrained models are the targets of SnatchML, which aims to hijack them for unrelated tasks involving different data distributions. Specifically, we hijack the models for: MNIST, CK+ for ER, CK+ for biometric identification, Olivetti for face recognition. We run SnatchML on the testset distribution to evaluate the hijacking task performance.
@@
For each hijacking attack, Figure 9 presents the performance under both black-box and white-box settings, along with the probability of a random guess. The results illustrate the capacity of overparametrised models to extract useful knowledge from a data distribution that is unrelated to their original training dataset. In fact, the SnatchML hijacking accuracy is comparable to the state-of-the-art accuracy of tasks like MNIST, when the victim model is trained on ImageNet.
VIII-B Cross modality: hijacking ECG for speech recognition
In the following, we investigate unrelated tasks where the original task involves a different modality than the hijacking task, i.e., hijacking ECG-based arrhythmia classification for speech recognition.
Setup. We evaluate three models pretrained on the MIT-BIH Arrhythmia Dataset: 1D-CNN, ResNet-18, and 1D-Inception. The adversary wants to hijack the models for speech recognition using the Free Spoken Digit (FSDD) Dataset [50]. The FSDD dataset consists of approximately $2,000$ audio recordings of spoken digits (0-9) by multiple speakers.
Results. The experimental results shown in Figure 10 highlight the effectiveness of SnatchML  in a cross-modality hijacking. These models achieved Top-1 accuracy around $23\%$ to $26\%$ and Top-5 accuracy ranging from $61\%$ to $64\%$.
Figure 10: SnatchML performance on hijacking an ECG models for speech classification on Free Spoken Digit dataset.
VIII-C Features visualization
To visualize the model’s capacity to distinguish classes from data distributions unrelated to the training set, in Figure 11, we present t-SNE plots comparing the feature distributions of MNIST digit classes using feature vectors extracted from a ResNet-18 model that is pretrained on ImageNet, and a randomly initialized ResNet-18. We observe distinct clusters corresponding to different digit classes in both cases, despite neither model being trained on the MNIST dataset. This observation confirms the previous results and illustrates the potential for model hijacking, even when the hijacking task is entirely unrelated to the original task.
Figure 11: Comparison of t-SNE distributions for MNIST digit classes using ResNet18 pretrained on ImageNet (left) and randomly initialized ResNet18 (right).
IX What is the impact of task complexity?
Figure 12: Effect of complexity ratio $r=m/n$ on the hijacking attack Accuracy.
Unlike existing work, our approach enables repurposing a model trained for a $n$-class classification task $\mathcal{T}$ for an $m$-class hijacking task $\mathcal{T^{\prime}}$, where $n\leq m$. In this section, we propose to investigate the extent of this property. Specifically, we explore the attack performance while varying the relative hijacking task complexity. For a given architecture, we use the number of classes as a proxy of the task complexity and define the complexity ratio $r=m/n$ as a metric to evaluate how difficult is the hijacking attack. The study was conducted using MobileNet, ResNet-9, 2D-CNN, and Transformer models, which were trained on MNIST. The number of training classes $n$ varied from only two digits to models trained to classify all $10$ digits in the complete dataset. Subsequently, the attack was performed to achieve facial recognition using Olivetti dataset, which contains $40$ identities. We conduct the attack with four different $m$ values by sampling different numbers of identities $m=40$, $m=20$, $m=8$, and $m=4$ identities. This resulted in 36 possible combinations, as shown in Figure 12. The scatter trend of Figure 12 reveals a clear inverse relationship between the ratio $m/n$ and the hijacking accuracy. As the ratio increases, the hijacking accuracy decreases. The number of classes in the hijacking task $m$ becomes significantly larger than the number of classes the model was originally trained on $n$, SnatchML’s ability to accurately hijack and classify the new classes diminishes. However, it is interesting to note that depending on the classes combination, SnatchML  achieved over $80\%$ Top-1 accuracy of the hijacking task, with $m=5\times n$ for ResNet-9, while related work do not work for $r>1$.
Figure 13: Correlation between the size of benign models trained on original task (ER on CK+ dataset) and their effectiveness in learning hijacking tasks: The results depict the variations in the accuracy of the hijacking tasks for user re-identification (First row) and user identification (Second row) when scaling the ML model size via model’s width expansion. The hijacked models are of different architectures (per column): 2D-CNN, ResNet-9, MobileNet, and Transformer.
X Why do models learn more than they should?
@@
In the previous sections, we show that SnatchML can hijack ML models to perform tasks that can be related or entirely unrelated to the original task. While the former is intuitive, the latter is more unexpected. To explore the underlying reasons, we examine two hypotheses: (i) Overparametrization of ML model and (ii) Random Projection.
X-A Correlation with model size
@@
ML models may learn features that are not particularly related to the original task. For example, in the context of ER, these models might capture facial characteristics that could inadvertently be used for biometric identification. From a conceptual perspective, such phenomenon might lead to critical security breaches, mainly for not respecting the fundamental principle of ”least privilege” in cybersecurity. Practically, the amount of parameters in a ML model reflects the upper bound of its learning capacity. The overparametrization implies that such models contain more parameters than necessary to accomplish the original task. In both software [46] and hardware [14] systems, undefined behavior and don’t-care states have been shown to be potential sources of vulnerabilities. In the same vein, we believe that this extra-capacity represents undefined behavior and don’t-care states with respect to the model that can be exploited by attackers. We suggest that the model’s capacity to unintentionally learn unnecessary features positively correlates with the accuracy of the hijacking task.
@@
To verify this hypothesis, we study the correlation between the ML model size and accuracy of the hijacking tasks for user re-identification and biometric identification. The results of our study are shown in Figure 13. We study the case of four ML models, notably 2D-CNN, ResNet-9, MobileNet, and Transformer. We vary the width expansion ratio (i.e., number of channels), from 0.25×0.25$\times$0.25 × to 2.5×2.5$\times$2.5 × to emulate different amounts of model parameters while maintaining the same baseline model architecture. From the reported results in Figure 13, we draw the following insights:
@@
Third, the observed trends of the model size and the accuracy of the hijacking task differ w.r.t the task complexity. For instance, ResNet-9 shows slight improvement of $2\%$ in the hijacking attack for user re-identification under overparameterized regime. For the identification task in Olivetti, ResNet-9 depicts higher improvement of $9\%$ in the hijacking attack ($2^{nd}$ row, $1^{st}$ and $2^{nd}$ columns in Figure 13). The re-identification task (i.e., the ER training dataset) includes $89$ user identities, making the re-identification task much more challenging than the identification task, which involves $\sim 40$ user identities.
X-B General case: Universality of Random DNNs
@@
The general case corresponds to a model that is hijacked for a task that is totally unrelated to the original task. The rationale by which features learnt by the model on the original task overlap with the hijacking task does not hold under these assumptions. In fact, as far as the hijacking task is concerned, this case is similar to a randomly initialised model. Our hypothesis is that the relatively high accuracy on the hijacking task can be explained by the universal approximation capabilities of random DNNs. We recall the Johnson-Lindenstrauss Lemma [24] which states that a set of points in a high-dimensional space can be projected into a lower-dimensional space while approximately preserving the pairwise distances between the points. While this lemma might give an interesting insight, it cannot rigorously explain the behavior of random ML models because of their non-linearity. However, several works such as [17, 4, 27, 37] investigated the capabilities of random neural networks. Particularly, Giryes et. al. [17] prove that DNN preserve the metric structure of the data as it propagates along the layers. Interestingly for our work, they show that networks tend to decrease the Euclidean distances between points with a small angle between them (“same class”), more than the distances between points with large angles between them (“different classes”). Further, the theoretical underpinnings provided by [17] demonstrate that deep random networks can act as powerful feature extractors, where the depth and width of the network play crucial roles in enhancing the representational power of the features.
Figure 14: Hijacking attack on a randomly initialized ResNet18 model to perform ER, Identification on CK+ and image classification on MNIST and CIFAR100.
@@
While the properties of random neural networks is extensively studied in the ML community, our work shows the potential security threats that can emanate from these properties. To illustrate this perspective, we run our hijacking attack on totally random neural networks, i.e., the parameters are sampled from a random Gaussian distribution. The results shown in Figure 14 are coherent with our observations in Section VIII. We provide a visualization of this setting in the latent space in Figure 11 in the Appendix.
XI Countermeasures
TABLE II: Model compression: accuracy of the original and hijacking tasks on different datasets and models.
[…]
Resnet-9 0.937 0.935 0.1$\times$ 0.257 0.217 0.293 0.294
Mobilenet 0.939 0.915 0.45$\times$ 0.148 0.143 0.176 0.163
Resnet-9 0.937 0.935 0.1$\times$ 0.232 0.144 0.479 0.359
Mobilenet 0.939 0.915 0.45$\times$ 0.074 0.070 0.076 0.066
Resnet-9 0.853 0.851 0.1$\times$ 0.764 0.541 0.787 0.756
Mobilenet 0.801 0.780 0.1$\times$ 0.544 0.482 0.751 0.633
XI-A Model Compression as a Defense
In Section X-A, we observed a correlation between the model size and the effectiveness of SnatchML. In this section, we explore to which extent model compression prevents SnatchML.
$f(.)=l^{n}\circ l^{n-1}\circ l^{n-2}\circ\dots\circ l^{2}\circ l^{1}s.t.\;\;f\in\mathcal{F}$ (3)
Where for each $\text{Where for each layer}\quad l^{j}=\{W_{1}^{j},W_{2}^{j},W_{3}^{j},...,W_{m_{j}}^{j}\}$ (4)
$f_{cmp}=\operatorname*{arg\,min}_{f_{m}\in\mathcal{F}}\alpha\cdot\mathcal{L}_{i}(f_{m}(X),Y)+\beta\cdot\mathcal{C}ount(f_{m})$ (5)
Setup. We design a search space of $14$ candidate width expansion ratios, ranging from 0.1×0.1$\times$0.1 × to 0.75×0.75$\times$0.75 ×. For each model $f$, we exhaustively iterate through the search space, sample an expansion ratio $m$, to get a compact model $f_{m}(.)$. At each iteration, we evaluate $f_{m}(.)$ using the loss $\mathcal{L}_{i}(.)$ (as a proxy for utility on the original task) and parameters count $\mathcal{C}ount(.)$ (as a proxy for model size). After the exhaustive search, we rank all candidate $f_{m}(.)$ according to their loss values and parameters count using the TOPSIS method [6] to render an optimal $f_{cmp}(.)$ with a tradeoff between utility and model size. We evaluate our compression method on three hijacking attacks: user re-identification on CK+ [28] identification on Olivetti [39], and Pneumonia recognition on Chest X-ray [25].
Results and Discussion. Results are detailed in Table II for each hijacking attack using three models: 2D-CNN, ResNet-9, and MobileNet. Across the $3$ datasets, compact models showed a notable reduction in hijacking accuracy. Specifically, ResNet-9 compact models demonstrate high resilience to hijacking while maintaining comparable accuracy to the baseline model. Identification accuracy on Olivetti and pneumonia type recognition on Chest X-ray decreases by $29\%$ and $38\%$, respectively, even under an aggressive compression strategy with a width expansion of $m=0.1\times$ on the baseline. Conversely, MobileNet shows a modest reduction in hijacking attack vulnerability for re-identification and identification tasks. This finding is consistent with the discussions in Section X-A, highlighting the intrinsic compactness of the baseline MobileNet model. Notably, in the pneumonia scenario, MobileNet with a $m=0.1\times$ shows low hijacking accuracy with a decrease of $11.3\%$ (Logits) and $15.7\%$ (FV).
It’s worth noting that in specific scenarios, e.g., re-identification on CK+, model compression does not always reduce the hijacking accuracy. This phenomenon may be attributed to the overlap between the original and hijacking tasks, particularly in smaller datasets where models are prone to overfitting. Such models tend to capture fine-grained features necessary for the original task, which concurrently exhibit a high correlation with the hijacking task. Isolating these fine-grained features could compromise the utility of the original task. Therefore, more sophisticated techniques of learning combined with model compression are needed.
XI-B Meta-unlearning:
In cases where the hijacking task can be be identified at training time, e.g., biometric identification, we propose to train the model such that it learns the original task and simultaneously unlearns the potentially malicious one. The intuition is that some internal representations are more transferable across tasks than others. For example, given an original task $\mathcal{T}_{i}$ and a related task $\mathcal{T}_{j}$, i.e., both $\mathcal{T}_{i},{T}_{j}\sim p(\mathcal{T})$, where $p(\mathcal{T})$ is distribution of tasks, a model might learn internal features that are broadly applicable to all tasks in $p(\mathcal{T})$, and others that are specific to the original task $\mathcal{T}_{i}$. Our objective is to penalize the learning of hijacking task-specific features while learning the ones relevant to the original one. The proposed approach is inspired from the meta-learning literature [15]: We train the model on $\mathcal{T}_{i}$, while maximising the loss function $\mathcal{L}_{j}$, relative to the (hijacking) task $\mathcal{T}_{j}$. The approach is detailed in Algorithm 1. Tables III and IV show the results of experiments designed to test the effectiveness of the meta-unlearning to retain accuracy on the original classification task under SnatchML. We also add a hijacking setting with stronger access than our approach, assuming the hijacking dataset can be used to train a neural network for hijacking (NN Surrogate), and not only using distances such as in SnatchML. Table IV shows the results for ’Gender classification’ as the original task hijacked for ’Ethnicity recognition’. It shows that when the model is trained with unlearning, the accuracy for the original task generally decreases. This is likely a consequence of the unlearning procedure, which may also remove features that are relevant to the original task. While the surrogate NN achieves, as expected, higher hijacking success, we interestingly notice that the meta-unlearning is more efficient on this attack than SnatchML, suggesting a higher robustness of distance-based inference in this setting. For example, the hijacking task accuracy (Ethnicity) has not been significantly impacted by the defense. These results were obtained with $\alpha=1$ and $\beta=0.01$ after empirical exploration. High unlearning coefficient leads to unacceptable accuracy drop in the original task.